### v1

@@ 2.3 The Hessian, Morse–Bott Functions and Asymptotic Behaviour (II)
As in the Euclidian case, linearizing the vector field $F$ at its equilibria sheds light on its local stability. Clearly, since $F=\operatorname{grad}f$ constitutes a gradient vector field, its linearization at an equilibrium $\mathbf{x}\in M$ is given by the Hessian $\operatorname{H}_{f}(\mathbf{x})$ of $f$ at $\mathbf{x}\in M$. In general, if $M$ is non-Euclidian, the computation of $\operatorname{H}_{f}(\mathbf{x})$ can be rather involved. Yet, at critical points $\mathbf{x}_{0}\in M$ of $f$, the Hessian is given by the symmetric bilinear form
@@
where $\sigma$ is any chart around $\mathbf{x}_{0}\in M$ and $\operatorname{\tilde{H}}(f\circ\sigma^{-1})$ denotes the ordinary Hesse matrix of the chart representation $f\circ\sigma^{-1}$. It is straightforward to show that (3) is independent of $\sigma$. Moreover, we will call $\mathbf{x}_{0}$ a strict saddle point if $\operatorname{H}_{f}(\mathbf{x}_{0})$ (or more precisely the associated symmetric operator) has at least one negative eigenvalue.
@@
Let $(M,g)$ be a Riemannian manifold and let $\mathbf{x}$ be a critical point of $f:M\to\mathbb{R}$. If $\operatorname{H}_{f}(\mathbf{x})$ is positive definite, then $\mathbf{x}$ is a strict local minimum of $f$.
Now the question arises whether the (asymptotic) stability of an equilibrium $\mathbf{x}\in M$ of (2) may dependent on the Riemannian metric $g$ — and the answer is surprisingly “yes”, cf. [57]. However, certain properties such as being a strict local minimum or an isolated critical point are obviously not up to the choice of the metric, and thus the (asymptotic) stability of those equilibria is independent of the Riemannian metric as the next theorem shows.
@@
  1. (a)
@@
Both assertions follow immediately from classical stability theory by taking $f$ as Lyapunov function, cf. [28, 29]. Handling non-isolated critical points is much more subtle. A first hint on how to approach this issue is obtained by Corollary 2.2 which could be restated (in a slightly weaker version) as follows:
@@
This suggests to work with Morse–Bott functions when it comes to non-isolated critical point. Recall that a smooth function $f:M\to\mathbb{R}$ is a called Morse–Bott function, when the set $C$ of critical points is a closed submanifolds of $M$ such that the tangent space of ${\sf T}_{\mathbf{x}}C$ coincides with the kernel of the Hessian operator for all $\mathbf{x}\in C$. Note that $C$ is allowed to have several connected components with possibly different dimensions, [28, p. 366], [48, Def. 2.41] or [5]. Thus Morse functions are particular Morse-Bott functions, where $C$ consists of 00-dimensional mainfolds (i.e. isolated points).
Now, the above concept allows a generalization of the Morse-Palais Lemma to Morse-Bott functions which is often called Morse-Bott Lemma, see [4]. It yields a local normal form of Morse-Bott functions near their critical points, which reads in local coordinates as follows
@@
with $\mathbf{x}\in\mathbb{R}^{n_{+}}$, $\mathbf{y}\in\mathbb{R}^{n_{-}}$, and $\mathbf{z}\in\mathbb{R}^{n_{0}}$, where $n_{+}$, $n_{-}$ and $n_{0}$ are the number of positive, negative and zero eigenvalues of the Hessian operator. This finally allows to prove that solutions of the respective gradient systems convergence to a single critical point.
Theorem 2.5.
Let $f:M\to\mathbb{R}$ be a Morse–Bott function on a Riemannian manifold $(M,g)$ with compact sublevel sets. Then every solution of the gradient flow (2) converges to a single critical point. Moreover for almost every initial point the flow converges to a local minimum.
Finally, we recall another very powerful result for analyzing the convergence of gradient systems which is based on Łojasiewicz’s celebrated gradient estimate [39]. Let $f:M\to\mathbb{R}$ be real analytic, ${\mathbf{x}}_{0}\in M$ a critical point of $f$ and assume w.l.o.g. $f({\mathbf{x}}_{0})=0$. Then near ${\mathbf{x}}_{0}\in M$ one has the estimate
$\|\operatorname{grad}(\mu\circ f)({\mathbf{x}})\|_{\mathbf{x}}\geq c\,,$ (4)
where $\mu:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a strictly increasing $C^{1}$-function and $c>0$ some positive constant, cf. [33, Cor. 1.1.25]. In the literature, one usually find $\mu(r):=r^{1-\theta}$ with $\theta\in(0,1)$. Eq. (4) easily allows to bound the length of any trajectory of (2) whose $\omega$-limit set is non-empty. Hence one gets the following result.
Theorem 2.6.
[39]. Let $(M,g)$ and $f:M\to\mathbb{R}$ be are real analytic. Then every non-empty $\omega$-limit set of (2) consists only of a single critical point.
@@ 2.4 The Exponential Map and Numerical Gradient Descent
Finally, we approach the problem of discretization of (2) resulting in a convergent gradient descent method. The ideas presented here can be traced back to [12, 13] and [55, 56]. Let
@@
denote the Riemannian exponential map at ${\mathbf{x}}\in M$, i.e. $t\to\exp_{\mathbf{x}}(t\xi)$ is the unique geodesic with initial value ${\mathbf{x}}\in M$ and initial velocity $\xi\in{\sf T}_{\mathbf{x}}M$. Here, we assume for simplicity that $(M,g)$ is (geodesically) complete, i.e. (5) is well-defined for the entire tangent space ${\sf T}_{\mathbf{x}}M$. Probably the simplest discretization scheme given by
@@
For compact Riemannian manifolds every trajectory of the discretized gradient descent (6) (with constant but small enough step size) converges to the set of critical points. Moreover, the set of initial points converging to strict saddle points has measure zero—in the sense of Sec. 2.1 here.
The second part of the above statement can be found in [37, Cor. 6] [^3] There the authors consider Riemannian manifolds embedded in the Euclidean space and retractions instead of an intrinsic Riemannian exponential function, but this does not restrict the generality of the result, cf. the Nash embedding theorem [7, Thm. 46]. Thus, for a function with compact sublevel sets, gradient descent (6) behaves similar as its continuous counter-part, cf. Cor. 2.2, i.e. it converges almost surely to a local minimum if the step size is chosen small enough.
Deeper results which yield convergence to a single critical point are more subtle to derive. Here we present one result in this direction which is again based on the analyticity of the cost function $f$ and on Łojasiewicz’s inequaliy.
@@
[33] If $(M,g)$ and $f$ are real analytic, and the step sizes are chosen according to a version of the first Wolfe–Powell condition for Riemannian manifolds, then pointwise convergence holds.
Remark 2.9.
It should be mentioned that gradient descent algorithms are usually studied as exact algorithms, not as numerical algorithms where real numbers are represented as floating point values and arithmetic is not exact. Numerical effects can be important, for instance cancellation effects from computing gradients in a naive way. Nevertheless, this paper will ignore these numerical issues.
Finally, in order to determine the largest admissible step size of our algorithm we need a notion of Lipschitz continuity of vector fields on Riemannian manifolds. Care has to be taken here since tangent vectors in different tangent spaces cannot be compared by default. Certainly, one can define local Lipschitz continuity via local charts but this does not allow to assign a meaningful Lipschitz constant to the vector field. A natural and intrinsic way to do this is to define Lipschitz continuity via parallel transport along the unique connecting geodesic, as is done in [24]. Here, we use the following equivalent approach via Riemannian normal coordinates which result from any orthogonal coordinate system on $T_{\mathbf{x}}M$ in combination with the Riemannian exponential map $\exp_{\mathbf{x}}$.
Definition 1.
Let $(M,g)$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be a smooth function. We say that $f$ is $\ell$-smooth around $\mathbf{x}\in M$ if there exist normal coordinates $\tilde{f}$ around $\mathbf{x}$ in which $\nabla\tilde{f}$ is $\ell$-Lipschitz.
@@ 3 Algorithm and Convergence
We will study the following randomly projected gradient descent algorithm. Given a Riemannian manifold $M$ of dimension $N$, a smooth function $f:M\to\mathbb{R}$, and a step size $\eta>0$ (with upper bound to be determined), the update rule is given by
$\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})),\quad g(\mathbf{x},\mathbf{u})=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle\,\mathbf{u},\quad\mathsf{S}_{\mathbf{x}_{i}}M\ni\mathbf{u}_{i}\overset{\mathrm{i.i.d}}{\sim}\mathcal{P}(\mathbf{x}_{i}),$ (7)
where $\mathsf{S}_{\mathbf{x}}M$ denotes the unit sphere in the tangent space $\mathsf{T}_{\mathbf{x}}M$ and $\mathcal{P}(\mathbf{x}_{i})$ some probability measure on the unit sphere $\mathsf{S}_{\mathbf{x}_{i}}M$. Intuitively, the gradient is projected onto a randomly chosen direction $\mathbf{u}$ at each step. Throughout we will consider two cases.
  • Either $\mathcal{P}(\mathbf{x}_{i})=\mathcal{U}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes the uniform (rotation invariant) probability distribution on the unit sphere, also called Haar measure;
  • or $\mathcal{P}(\mathbf{x}_{i})=\mathcal{D}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes a finite probability distribution on the unit sphere in some sense approximating a uniform distribution.
Let us clarify the case of a finite probability distribution. In this case we have $k$ continuous vector fields $\xi_{j}$ on $M$ and continuous weights $p_{j}:M\to\mathbb{R}$ with $p_{j}>0$ and $\sum_{j_{1}}^{k}p_{j}\equiv 1$. Then $\mathbf{u}_{i}$ takes value $\xi_{j}(\mathbf{x}_{i})$ with probability $p_{j}(\mathbf{x}_{i})$. Moreover we assume that at every point $\mathbf{x}$, the the tangent vectors $\xi_{1}(\mathbf{x}),\dots,\xi_{k}(\mathbf{x})$ span the entire tangent space even after removing all vectors collinear with some given direction $\xi_{j}(\mathbf{x})$.
Our goal is to analyze the convergence behaviour of this algorithm, and in particular to show that it converges almost surely to a local minimum of $f$. A deterministic version of the above described algorithm was analysed in [1, Def. 4.2.1, Thm. 4.3.1] under the heading “gradient-related methods”.

### v2

@@ 2.3 The Hessian, Morse–Bott Functions and Asymptotic Behaviour (II)
As in the Euclidian case, linearizing the vector field $F$ at its equilibria sheds light on its local stability. Clearly, since $F=-\operatorname{grad}f$ constitutes a gradient vector field, its linearization at an equilibrium $\mathbf{x}\in M$ is given by the Hessian $\operatorname{H}_{f}(\mathbf{x})$ of $f$ at $\mathbf{x}\in M$. In general, if $M$ is non-Euclidian, the computation of $\operatorname{H}_{f}(\mathbf{x})$ can be rather involved. Yet, at critical points $\mathbf{x}_{0}\in M$ of $f$, the Hessian is given by the symmetric bilinear form
@@
where $\sigma:U\to\mathbb{R}^{n}$ is any chart around $\mathbf{x}_{0}\in M$ and $\operatorname{\tilde{H}}(f\circ\sigma^{-1})$ denotes the ordinary Hesse matrix of the chart representation $f\circ\sigma^{-1}$. It is straightforward to show that (3) is independent of $\sigma$. Moreover, we will call $\mathbf{x}_{0}$ a strict saddle point if $\operatorname{H}_{f}(\mathbf{x}_{0})$ (or more precisely the associated symmetric operator) has at least one negative eigenvalue.
@@
Let $M$ be a Riemannian manifold and let $\mathbf{x}$ be a critical point of the smooth function $f:M\to\mathbb{R}$. If $\operatorname{H}_{f}(\mathbf{x})$ is positive definite, then $\mathbf{x}$ is a strict local minimum of $f$.
Now the question arises whether the (asymptotic) stability of an equilibrium $\mathbf{x}\in M$ of (2) may dependent on the Riemannian metric $\mathsf{g}$ — and the answer is surprisingly “yes”, cf. [67]. However, certain properties such as being a strict local minimum or an isolated critical point are obviously not up to the choice of the metric, and thus the (asymptotic) stability of those equilibria is independent of the Riemannian metric as the next theorem shows.
@@
Let $M$ be a Riemannian manifold and $f:M\to\mathbb{R}$ a smooth function. Then the following hold:
  1. (i)
@@
Both assertions follow immediately from classical stability theory by taking $f$ as Lyapunov function, cf. [33, 34]. Handling non-isolated critical points is much more subtle. A first hint on how to approach this issue is obtained by Cor. 2.2 which could be restated (in a slightly weaker version) as follows:
@@
This suggests to work with Morse–Bott functions when it comes to non-isolated critical points. Recall that a smooth function $f:M\to\mathbb{R}$ is a called Morse–Bott function, if the set $C$ of critical points is a closed submanifold of $M$, such that its tangent space ${\sf T}_{\mathbf{x}}C$ coincides with the kernel of the Hessian operator for all $\mathbf{x}\in C$. Note that $C$ is allowed to have several connected components with possibly different dimensions, [33, p. 366], [57, Def. 2.41] or [7]. Thus Morse functions are particular Morse–Bott functions, where $C$ consists only of 00-dimensional mainfolds (i.e. isolated points).
Now, the above concept allows a generalization of the Morse–Palais Lemma to Morse–Bott functions which is often called Morse–Bott Lemma, see [6]. It yields a local normal form of Morse–Bott functions near their critical points, which reads in local coordinates as follows
@@
with $\mathbf{x}\in\mathbb{R}^{n_{+}}$, $\mathbf{y}\in\mathbb{R}^{n_{-}}$, and $\mathbf{z}\in\mathbb{R}^{n_{0}}$, where $n_{+}$, $n_{-}$ and $n_{0}$ are the numbers of positive, negative and zero eigenvalues of the Hessian operator respectively. Here $\|\cdot\|^{2}$ denotes the usual Euclidean norm. This finally allows to prove that solutions of the respective gradient system converge to a single critical point.
Theorem 2.5 (Thm. 2.3 in [4]).
Let $f:M\to\mathbb{R}$ be a Morse–Bott function on a Riemannian manifold $M$ with compact sublevel sets. Then every solution of the gradient flow (2) converges to a single critical point. Moreover, for almost every initial point, the flow converges to a local minimum.
Finally, we recall another very powerful result for analyzing the convergence of gradient systems which is based on Łojasiewicz’s celebrated gradient estimate [46]. Let $f:M\to\mathbb{R}$ be real analytic, ${\mathbf{x}}_{0}\in M$ a critical point of $f$, and assume without loss of generality that $f({\mathbf{x}}_{0})=0$. Then, near ${\mathbf{x}}_{0}$, one has the estimate
$\|\operatorname{grad}(\mu\circ f)({\mathbf{x}})\|\geq c\,,$ (4)
where $\mu:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a strictly increasing $C^{1}$-function and $c>0$ some positive constant, cf. [39, Cor. 1.1.25]. In the literature, one usually finds $\mu(r):=r^{1-\theta}$ with $\theta\in(0,1)$. Eq. (4) easily allows to bound the length of any trajectory of (2) whose $\omega$-limit set is non-empty. Hence one gets the following result.
Theorem 2.6 ([46]).
Let $(M,\mathrm{g})$ and $f:M\to\mathbb{R}$ be are real analytic. Then every non-empty $\omega$-limit set of (2) consists only of a single critical point.
@@ 2.4 The Exponential Map and Numerical Gradient Descent
Finally, we approach the problem of discretization of (2) resulting in a convergent gradient descent method (cp. Fig.1). The ideas presented here can be traced back to [15, 16] and [64, 65]. Let
@@
denote the Riemannian exponential map at ${\mathbf{x}}\in M$, i.e. $t\to\exp_{\mathbf{x}}(t\xi)$ is the unique geodesic with initial value ${\mathbf{x}}\in M$ and initial velocity $\xi\in{\sf T}_{\mathbf{x}}M$. Here, we assume for simplicity that $(M,\mathsf{g})$ is (geodesically) complete, i.e. (5) is well-defined for the entire tangent space ${\sf T}_{\mathbf{x}}M$. Probably the simplest discretization scheme given by
@@
Remark 2.7.
It should be mentioned that gradient descent algorithms are usually studied as exact algorithms, not as numerical algorithms where real numbers are represented as floating point values and arithmetic is not exact. Numerical effects can be important, for instance cancellation effects from computing gradients in a naive way [54]. Nevertheless, this paper will ignore these numerical issues.
@@
If $f$ has compact sublevel sets, every trajetory of the discretized gradient descent (6) (with constant but small enough step size) converges to the set of critical points. Moreover, if additionally
  1. (i)
    all critical points are isolated, then any trajectory of (6) converges to a single critical point of $f$.
  2. (ii)
    all saddle points are strict, then the set of initial points converging to strict saddle points has measure zero (in the sense of Sec. 2.1).
Part (ii) of the above statement can be found in [44, Cor. 6]. [^7] There the authors consider Riemannian manifolds embedded in the Euclidean space and retractions instead of an intrinsic Riemannian exponential function, but this does not restrict the generality of the result, cf. the Nash embedding theorem [9, Thm. 46]. Thus, for a function with compact sublevel sets, gradient descent (6) behaves similarly to its continuous counter-part, cf. Prop. 2.1 and Cor. 2.2, i.e. it converges almost surely to a local minimum if the step size is chosen small enough.
Deeper results, which yield convergence to a single critical point, are more subtle to derive. Here we present one result in this direction which is again based on the analyticity of the cost function $f$ and on Łojasiewicz’s inequaliy.
@@
[39] If $(M,\mathsf{g})$ and $f$ are real analytic, and the step sizes are chosen according to a version of the first Wolfe–Powell condition for Riemannian manifolds, then pointwise convergence holds.
Finally, in order to determine the largest admissible step size of our algorithm, we need a notion of Lipschitz continuity of vector fields on Riemannian manifolds. Care has to be taken here since tangent vectors in different tangent spaces cannot be compared by default. Certainly, one can define local Lipschitz continuity via local charts but this does not allow to assign a meaningful Lipschitz constant to the vector field. A natural and intrinsic way to do this is to define Lipschitz continuity via parallel transport along the unique connecting geodesic, as is done in [28].
Definition 2.10.
Let $(M,\mathsf{g})$ be a complete Riemannian manifold and let $X$ be a vector field on $M$. We say that $X$ is $\ell$-Lipschitz if for all $\mathbf{x},\mathbf{y}\in M$ with [^8] Here $r_{M}$ stands for the injectivity radius of $M$. $d(x,y)\leq r_{M}$ it holds that [^9] Here $\Pi_{\mathbf{y},\mathbf{x}}$ denotes the parallel transport from $\mathbf{y}$ to $\mathbf{x}$ along the unique length minimizing geodesic connecting the two points. $\|X(\mathbf{x})-\Pi_{\mathbf{y},\mathbf{x}}X(\mathbf{y})\|_{\mathbf{x}}\leq\ell\,d(\mathbf{x},\mathbf{y})$. If $f:M\to\mathbb{R}$ is a differentiable function, then we say that $f$ is $\ell$-smooth if $\operatorname{grad}(f)$ is $\ell$-Lipschitz.
Note that this definition is slightly broader than the definition given in [28].
@@ 3 Algorithm and Convergence
We will study the following randomly projected gradient descent algorithm (cp. Fig. 1). Given a Riemannian manifold $M$ of dimension $N$, an initial state $\mathbf{x}_{0}\in M$, a smooth function $f:M\to\mathbb{R}$, and a step size $\eta>0$ (with upper bound to be determined), the update rule is given by
$\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})),\,\,\,g(\mathbf{x},\mathbf{u})=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle\,\mathbf{u},\,\,\,\mathsf{S}_{\mathbf{x}_{i}}M\ni\mathbf{u}_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{P}(\mathbf{x}_{i}),$ (7)
where $\mathsf{S}_{\mathbf{x}}M$ denotes the unit sphere in the tangent space $\mathsf{T}_{\mathbf{x}}M$ and $\mathcal{P}(\mathbf{x}_{i})$ some probability distribution on the unit sphere $\mathsf{S}_{\mathbf{x}_{i}}M$. (Recall that “i.i.d.” stands for “independent and identically distributed”). Intuitively, the gradient is projected onto a randomly chosen direction $\mathbf{u}$ at each step. Throughout this paper we will consider two cases:
  1. (I)
    Either $\mathcal{P}(\mathbf{x}_{i})=\mathcal{U}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes the uniform (rotationally invariant) probability distribution on the unit sphere, also called Haar measure;
  2. (II)
    or $\mathcal{P}(\mathbf{x}_{i})=\mathcal{D}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes a finite probability distribution on the unit sphere accessing all possible directions in the sense of Assumption 3.1.
Assumption 3.1 (Discrete case).
Given $k$ continuous vector fields $\xi_{1},\ldots,\xi_{k}$ and $k$ continuous weights $p_{1},\ldots,p_{k}$ on $M$ with $p_{j}\geq 0$ and $\sum_{j=1}^{k}p_{j}\equiv 1$, such that $p_{j}\xi_{j}/\|\xi_{j}\|$ is continuous, where the expression is set to zero when it is undefined. This implies for instance that $p_{j}(\mathbf{x})=0$ whenever $\xi_{j}(\mathbf{x})=0$.
Then $\mathbf{u}_{i}$ takes value $\xi_{j}(\mathbf{x}_{i})/\|\xi_{j}(\mathbf{x}_{i})\|$ with probability $p_{j}(\mathbf{x}_{i})$. Moreover, we suppose that at every point $\mathbf{x}$, the tangent vectors $p_{1}(\mathbf{x})\xi_{1}(\mathbf{x}),\dots,p_{k}(\mathbf{x})\xi_{k}(\mathbf{x})$ span the entire tangent space even after removing any single vector $p_{j}(\mathbf{x})\xi_{j}(\mathbf{x})$ and, if present, all its scalar multiples.
The following technical lemma will be useful later.
Lemma 3.2.
Let $K$ be a compact subset of a manifold $M$, and let $\{v_{j}\}_{j=1}^{k}$ be a set of continuous vector fields on $K$ which span the tangent space at every point. Then, for any continuous, non-vanishing vector field $v$ on $K$ there is some $\varepsilon>0$ such that
${\min_{\mathbf{x}\in K}\max_{j=1,\ldots,k}|\langle v_{j}(\mathbf{x}),v(\mathbf{x})\rangle|=\varepsilon.}$
Proof.
The functions $\mathbf{x}\mapsto|\langle v_{j}(\mathbf{x}),v(\mathbf{x})\rangle|$ defined on $K$ are continuous and non-negative. Hence their maximum over $j$ is still a continuous non-negative function on $K$. Since the $v_{j}$ span each tangent space, this function is even strictly positive, and by compactness of $K$, there is a positive global minimum. ∎
Note that a related condition is used in [32, Thm. 5.2] to guarantee convergence to the set of critical points.
Our goal is to analyze the convergence behavior of this algorithm, and in particular, to show that it converges almost surely to a local minimum of $f$. A deterministic version of the above described algorithm (in the sense that all “perturbations” from the classical gradient descent are deterministic and not random) was analyzed in [1, Def. 4.2.1, Thm. 4.3.1] under the heading “gradient-related methods”.